ASTRONOMICAL AND ATMOSPHERIC SPECTROSCOPY IN THE SMM/THz: EXPERIMENTS, ANAL-YSIS, AND CATALOGS.

\underline{FRANK C. DE LUCIA}, Department of Physics, 191 W. Woodruff Ave., Ohio State University, Columbus, OH 43210 USA.

For many years the needs of the astronomical and atmospheric science communities have directly or indirectly motivated much of the work at this symposium. Interestingly, the power and sophistication of the field instruments have grown at such a rate that the needs for laboratory data seem to be diverging rather than converging. The central role played by catalogs will be discussed, as well as their impact on the spectroscopic community.